OpenStudy (anonymous):

This question was on SAT, and no joke, it needs clarification, to me it looks just invalid!! 3. If x and y are integers, and 3x + 2y = 13, which of the following could be the value of y ?

OpenStudy (anonymous):

\[0\]\[1\]\[2\]\[3\]\[4\]

OpenStudy (anonymous):

I know it is no zero, right?

OpenStudy (jdoe0001):

I'd solve for "y" first, and then plug in the values and solve for "x", see if which one gives you an integer

OpenStudy (anonymous):

\[y=-2/3x+13/3\] @jdoe0001, ty, I think if I think, I'll get it...

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

b/c for other values it is not an integer.

OpenStudy (anonymous):

TY!

OpenStudy (jdoe0001):

\(\bf 3x+2y=13\implies y = \cfrac{13-3x}{2}\)

OpenStudy (jdoe0001):

so... you'd test say... 0 , if y = 0, then \(\bf 0 = \cfrac{13-3x}{2}\implies x=\cfrac{13}{3} \) so no dice on 0

OpenStudy (jdoe0001):

so checking on y = 1 then ... \(\bf 1 = \cfrac{13-3x}{2}\implies 3x+2=13\implies x=\cfrac{11}{3}\) so no dice on 1 either

OpenStudy (jdoe0001):

and so on :)

OpenStudy (anonymous):

So this SAT question would take at least about 6 minutes. That's not a good time per question on the SAT.

OpenStudy (amistre64):

instead of solving for y; solve for x .. they give you options in terms of y

OpenStudy (amistre64):

3x + 2y = 13 \[x = \frac{-2y+13}{3}\] which value of y makes 13-2y a multiple of 3?

OpenStudy (amistre64):

13 11 9 7 5

OpenStudy (amistre64):

takes less then 30 seconds for me